[6:28 pm, 17/02/2022] Anshuman✌️: Suppose that business travelers and vacationers have

the following demand for airline tickets from New York

to Boston:

QUANTITY DEMANDED QUANTITY DEMANDED

PRICE (BUSINESS TRAVELERS) (VACATIONERS)

$150 2,100 1,000

200 2,000 800

250 1,900 600

300 1,800 400

a. As the price of tickets rises from $200 to $250, what

is the price elasticity of demand for (i) business

travelers and (ii) vacationers? (Use the midpoint

method in your calculations.)

b. Why might vacationers have a different elasticity

than business travelers?


1
Expert's answer
2022-02-21T13:09:38-0500

"PED= \frac{dQ}{Q}/\frac{dP}{P}"

Where:

"PE" - price elasticity

"Q" - Quantity f the demanded good

"P" - Price of the demanded good



(I) Price elasticity of demand for business travelers


Price elasticity of demand "=\frac{1,900 -2,000}{1,900+2,000}/\frac {250-200}{250+200}"



Price elasticity of demand "= \frac{-100}{3,900}/\frac{50}{450} = -0.2307"


"PED= 0.2307"


(ii) Price elasticity of demand for vacation travelers


Price elasticity of demand ="=\frac{600-800}{600+800} /\frac{250-200}{250+200}"



Price elasticity of demand ="= \frac{-200}{1400}/\frac{50}{450} = -1.2867"


"PED =1.2867"


b)The price elasticity of demand for vacationers is higher than the elasticity for business travelers because vacationers can choose more easily a different mode of transportation (like driving or taking the train) or not travel at all.